OpenStudy (anonymous):

integrate from 0 to 2 (squareroot of (6x+4)-2x

OpenStudy (anonymous):

\[\int\limits_0^2 \sqrt{6x+4}dx -2 \int\limits_0^2 x dx\] Let \[\xi=6x+4 \implies d \xi=6 dx \implies dx=\frac{d \xi}{6}\] \[\frac{1}{6} \int\limits_4^{16}\sqrt{\xi}d \xi-2 \int\limits_0^2 x dx\]

OpenStudy (anonymous):

ty!

OpenStudy (anonymous):

wait im kinda confused uhm wat i did was

OpenStudy (anonymous):